The sitting height (from seat to top of head) of drivers must be considered in the design of a new car model. Men have sitting heights that are normally distributed with a mean of 36.0 inches and a standard deviation of 1.4 inches. Engineers have provided plans that can accommodate men with sitting heights up to 38.8 heights, but taller men cannot fit. If a man is randomly selected, find the probability that he has a sitting height less than 38.8 inches.
Is it possible for a customer paying by check to spend more than $185? Explain.

Answer by stanbon(57347)

The sitting height (from seat to top of head) of drivers must be considered in the design of a new car model. Men have sitting heights that are normally distributed with a mean of 36.0 inches and a standard deviation of 1.4 inches. Engineers have provided plans that can accommodate men with sitting heights up to 38.8 heights, but taller men cannot fit. If a man is randomly selected, find the probability that he has a sitting height less than 38.8 inches.
---
z(38.8) = (38.8-36)/1.4 = 2
----
P(x < 38.8) = P(z < 2) = normalcdf(-100,2) = 0.9772
=====================
Cheers,
Stan H.
=====================